# Cheap Learning: Maximising Performance of Language Models for Social Data Science Using Minimal Data

### v1

@@ 5 Conclusions
Second, perhaps most of all, it was striking how zero-shot prompt engineering performed well across all of the tasks we reviewed, comfortably outperforming any of the other methods on the movie review task and offering equivalent performance on the abuse classification task. Considering these techniques require no labelled training data (though still require test data) and are very easy to implement, they appear to offer a very promising and flexible solution to automatic content analysis in the future. Their performance on a wider variety of social science concepts needs to be tested; and it should also be borne in mind that there is a financial cost to using them (though not one that will be significant unless millions of lengthy text documents are being labelled). Furthermore, larger models than the current generation of LLMs are already in development, and it is reasonable to expect both performance gains and price decreases as they emerge (in all of our experiments, the increase in size from GPT-3 to GPT-4 corresponded with an increase in performance). As this next generation emerges, it may be that other methods of automatic content analysis will start to become redundant, though this proposition would need to be rigorously tested.
Lastly, despite the impressive performance of zero-shot prompt engineering using LLMs, it is important to assess LLMs across various types of tasks to understand their capacity and limitations. For instance, our results show that model performance can vary across seeds or the prompts tested. This instability further highlights the importance of robustness and reproducibility in LLMs. Furthermore, LLMs can be biased towards certain demographics or identities when, for example, asked to perform curriculum screening (Dastin, 2022; Veldanda et al., 2023). In this paper, we solely consider binary classification tasks. How LLMs perform on more complicated tasks that involve language understanding such as logical reasoning and inference remains an open research question.
@@ Acknowledgments.
This work was supported by the Public Policy Programme of The Alan Turing Institute under the EPSRC grant EP/N510129/1. This work was supported by the Ecosystem Leadership Award under the EPSRC Grant EP/X03870X/1 and The Alan Turing Institute. Funded by the UK Government
@@ A.1.1 Transfer learning
In Figure 5 we present the learning curves for transfer learning experiments under balanced train and balanced test (5 a), unbalanced train and unbalanced test (5 b), and balanced train and unbalanced test (5 c) using DeBERTa-v3 and DistilBERT. Both models yield similar performance across labelling budgets, while a small fluctuation is observed at 256 training points. With 512 training points models obtain (close to) best results in terms of F1 score, suggesting that 512 can be the amount of data required for reaching optimal performance using transfer learning. When trained on balanced data, both models show a striking jump in performance in the early stage of training with 128 training training points.
@@ A.1.2 Prompt engineering
In Figure 6 we present the learning curves for prompt engineering experiments under balanced train and balanced test (6 a), unbalanced train and unbalanced test (6 b), and balanced train and unbalanced test (6 c) using GPT-2 and DistilBERT. The results are based on the prompt ‘Does this text contain abuse?’. Generally, GPT-2 obtains better results than DistilBert, which can be due to model structures and pre-training objectives. Similar to transfer learning, models trained under balanced scenarios demonstrate a faster learning curve at distinguishing between classes in comparison with unbalanced settings. Best results are also obtained at 512 training points. In fact, the classification performance of both models drops at the maximum labelling budgets under balanced training scenarios, which shows that more data is not always beneficial for model learning. This emphasizes the advantages of class balance outweigh data size. As reported in previous work [Islam et al., 2017], high variance across seeds is of particular interest and highlights the need for searching for the best seed.
@@
To investigate the impact of different prompts on classification, in Figure 7 we plot the Macro F1 for each of the three prompts under under balanced train and balanced test (7 a), unbalanced train and unbalanced test (7 b), and balanced train and unbalanced test (7 c). The results shown are based on Distilbert. Model performance expectedly varies across prompts while demonstrating similar learning curves over labelling budgets. The confidence intervals of the learning curves of each run overlap, which suggests that divergences across seeds can be more important than the potentials of prompt types. This finding is similar to the work done by Le Scao and Rush [2021]. Best results achieved are with ‘Does this text contain abuse?’. The reasons that led to the differences in prompts are unclear. Future work can explore prompt types to improve the classification by changing, for instance, the syntactical, lexical and semantic aspects of prompts.
@@ A.2 Binary movie sentiment classification task
Model Number of points
Prompt
Macro F1 score
GPT-3 855
Using one word, does the movie review contain negative sentiment, Yes or No?
0.67
GPT-3 855
Using one word, classify the sentiment of the movie review using ’Positive’ or ’Negative’.
0.85
GPT-3 854
You are a researcher who needs to classify movie reviews as containing negative sentiment or not containing negative sentiment. Using one word, does the movie review contain negative sentiment, Yes or No?
0.11
GPT-3.5 852
Using one word, does the movie review contain negative sentiment, Yes or No?
0.92
GPT-3.5 843
Using one word, classify the sentiment of the movie review using ’Positive’ or ’Negative’.
0.90
GPT-3.5 851
You are a researcher who needs to classify movie reviews as containing negative sentiment or not containing negative sentiment. Using one word, does the movie review contain negative sentiment, Yes or No?
0.90
GPT-4 855
Using one word, does the movie review contain negative sentiment, Yes or No?
0.93
GPT-4 855
Using one word, classify the sentiment of the movie review using ’Positive’ or ’Negative’.
0.94
GPT-4 855
You are a researcher who needs to classify movie reviews as containing negative sentiment or not containing negative sentiment. Using one word, does the movie review contain negative sentiment, Yes or No?
0.88
@@
A.3 Labelling Functions
@@
Task
Type Label
Explanation
Coverage Overalap
Binary abuse
Keywords ABUSE
If any of the words considered abusive from a given list of keywords is found in the text, label the text as abusive.
0.36 0.36
Binary abuse
Heuristic ABUSE
If the length of the text is above 9500 characters (limit of 10 000 characters), label the text as abusive.
0.00 0.00
Binary abuse
NLP ABUSE
If the polarity of the text is below -0.25, label the text as abusive.
0.07 0.07
Binary abuse
NLP NOT_ABUSE
If the polarity of the text is equal or more than 0.1, label the text as non abusive.
0.41 0.41
Binary abuse
Annotators Both
Take the annotators’ data and use it as labelling functions: given an annotator $i$ labelling a set of texts, the annotator-function labels the previously annotated texts and abstains with all non-annotated ones.
0.02 0.02
Binary movie sentiment
Keywords Positive
If any of the words considered related to positive sentiments is found from a given list of keywords, label the text as positive.
0.90 0.90
Binary movie sentiment
Keywords Negative
If any of the words considered related to negative sentiments is found from a given list of keywords, label the text as negative.
0.62 0.62
Binary movie sentiment
Regex Positive
If any expression related to the acting, the direction, the edition, the sound or in general to the film or the movie contains a positive keyword, then label the text as positive.
0.25 0.25
Binary movie sentiment
Regex Negative
If any expression related to the acting, the direction, the edition, the sound or in general to the film or the movie contains a negative keyword, then label the text as negative.
0.03 0.03
Binary movie sentiment
NLP Negative
If the polarity of the text is below -0.05, label it as negative.
0.03 0.03
Binary movie sentiment
NLP Negative
If the polarity of the text is above 0, label it as positive.
0.95 0.95
Binary movie sentiment
NLP Positive
If the subjectivity of the text is below 0.3, label it as positive.
0.01 0.01
@@
A.4 Zero-shot Prompt Engineering
@@
Task
Prompt
GPT-3 GPT-3.5 GPT-4
N Non-response
Binary abuse
Using one word, does the internet comment contain toxic language, Yes or No?
0.721 0.733 0.781
21
Binary abuse
Using one word, is this internet comment using toxic language, Yes or No?
0.667 0.684 0.797
13
Binary abuse
You are a researcher who needs to classify comments on the internet as containing abusive language or not containing abusive language. Using one word, does the internet comment contain abusive language, Yes or No?
0.713 0.802 0.813
19
Binary movie sentiment
Using one word, does the movie review contain negative sentiment, Yes or No?
0.952 0.923 0.961
2
Binary movie sentiment
Using one word, classify the sentiment of the movie review using ’Positive’ or ’Negative’.
0.941 0.945 0.964
8
Binary movie sentiment
You are a researcher who needs to classify movie reviews as containing negative sentiment or not containing negative sentiment. Using one word, does the movie review contain negative sentiment, Yes or No?
0.876 0.918 0.939
7
@@
Task
Non-response
Binary abuse
Cannot determine if the internet comment contains abusive language or not based on the given text
Binary abuse
I cannot determine if this internet comment contains abusive language or not as it is not a complete sentence
Binary abuse
The internet comment does not contain a clear indication of abusive language
Binary abuse
The given internet comment does not contain any language
Binary abuse
I cannot determine if the internet comment contains abusive language or not as the comment is incomplete and ends abruptly
Binary movie sentiment
Neutral
Binary movie sentiment
Mixed

### v2

@@ 5 Conclusions
Second, perhaps most of all, it was striking how zero-shot prompt engineering performed well across all of the tasks we reviewed, comfortably outperforming any of the other methods on the movie review task and offering equivalent performance on the abuse classification task. However, even if their performances are high, over-predictions and thus systematic biases can also be present. An analysis of these elements are needed no matter the technique used to avoid arriving at incorrect conclusions. Considering these techniques require no labelled training data (though still require test data) and are very easy to implement, they appear to offer a very promising and flexible solution to automatic content analysis in the future. Their performance and biases on a wider variety of social science concepts need to be tested; and it should also be borne in mind that there is a financial cost to using them (though not one that will be significant unless millions of lengthy text documents are being labelled). Furthermore, larger models than the current generation of LLMs are already in development, and it is reasonable to expect both performance gains and price decreases as they emerge (in all of our experiments, the increase in size from GPT-3 to GPT-4 corresponded with an increase in performance and a decrease in over-prediction). As this next generation emerges, it may be that other methods of automatic content analysis will start to become redundant, though this proposition would need to be rigorously tested. This proposition would be of particular interest as open- and closed-source LLMs can present serious biases that could skew their uses (Spirling, 2023; Ashwin et al., 2023).
Lastly, despite the impressive performance of zero-shot prompt engineering using LLMs, it is important to assess LLMs across various types of tasks to understand their capacity and limitations. For instance, our results show that model performance and over-predictions can vary across seeds or the prompts tested. This instability further highlights the importance of robustness and reproducibility in LLMs. Furthermore, LLMs can be biased towards certain demographics or identities when, for example, asked to perform curriculum screening (Dastin, 2022; Veldanda et al., 2023), or annotating refugees’ interviews (Ashwin et al., 2023). In this paper, we solely consider binary classification tasks as our intention is to keep results as relatable as possible so that the reader can understand the concepts and bring this knowledge to their classification tasks, as complex as they could be. While an open question remains on how LLMs perform on more complicated tasks that involve language understanding such as logical reasoning and inference, we contribute by giving social scientists different strategies to tackle challenges where data scarcity and bias might be present like analysing interview data from under-represented groups (Deterding and Waters, 2021; Small and Calarco, 2022; Bonikowski and Nelson, 2022; Ashwin et al., 2023; Atari et al., 2023) or text analysis and labelling of historical archives (Jaillant, 2022; Marciano et al., 2018; Rahal et al., 2024).
@@ Acknowledgments.
This work was supported by the Public Policy Programme of The Alan Turing Institute under the EPSRC grant EP/N510129/1. This work was supported by the Ecosystem Leadership Award under the EPSRC Grant EP/X03870X/1 and The Alan Turing Institute. Funded by the UK Government.
@@
Conflict of Interest Statement.
The authors declared no potential conflicts of interest with respect to the research, authorship, and publication of this article.
@@ A.1.1 Transfer learning
In Figure 7 we present the learning curves for transfer learning experiments under balanced train and balanced test (7 a), unbalanced train and unbalanced test (7 b), and balanced train and unbalanced test (7 c) using DeBERTa-v3 and DistilBERT. Both models yield similar performance across labelling budgets, while a small fluctuation is observed at 256 training points. With 512 training points models obtain (close to) best results in terms of F1 score, suggesting that 512 can be the amount of data required for reaching optimal performance using transfer learning. When trained on balanced data, both models show a striking jump in performance in the early stage of training with 128 training training points.
@@ A.1.2 Prompt engineering
In Figure 8 we present the learning curves for prompt engineering experiments under balanced train and balanced test (8 a), unbalanced train and unbalanced test (8 b), and balanced train and unbalanced test (8 c) using GPT-2 and DistilBERT. The results are based on the prompt ‘Does this text contain abuse?’. Generally, GPT-2 obtains better results than DistilBert, which can be due to model structures and pre-training objectives. Similar to transfer learning, models trained under balanced scenarios demonstrate a faster learning curve at distinguishing between classes in comparison with unbalanced settings. Best results are also obtained at 512 training points. In fact, the classification performance of both models drops at the maximum labelling budgets under balanced training scenarios, which shows that more data is not always beneficial for model learning. This emphasizes the advantages of class balance outweigh data size. As reported in previous work [Islam et al., 2017], high variance across seeds is of particular interest and highlights the need for searching for the best seed.
@@
To investigate the impact of different prompts on classification, in Figure 9 we plot the Macro F1 for each of the three prompts under under balanced train and balanced test (9 a), unbalanced train and unbalanced test (9 b), and balanced train and unbalanced test (9 c). The results shown are based on Distilbert. Model performance expectedly varies across prompts while demonstrating similar learning curves over labelling budgets. The confidence intervals of the learning curves of each run overlap, which suggests that divergences across seeds can be more important than the potentials of prompt types. This finding is similar to the work done by Le Scao and Rush [2021]. Best results achieved are with ‘Does this text contain abuse?’. The reasons that led to the differences in prompts are unclear. Future work can explore prompt types to improve the classification by changing, for instance, the syntactical, lexical and semantic aspects of prompts.
@@ A.2 Binary movie sentiment classification task
Model Number of points Prompt Macro F1 score
GPT-3 855 Using one word, does the movie review contain negative sentiment, Yes or No? 0.67
GPT-3 855 Using one word, classify the sentiment of the movie review using ’Positive’ or ’Negative’. 0.85
GPT-3 854 You are a researcher who needs to classify movie reviews as containing negative sentiment or not containing negative sentiment. Using one word, does the movie review contain negative sentiment, Yes or No? 0.11
GPT-3.5 852 Using one word, does the movie review contain negative sentiment, Yes or No? 0.92
GPT-3.5 843 Using one word, classify the sentiment of the movie review using ’Positive’ or ’Negative’. 0.90
GPT-3.5 851 You are a researcher who needs to classify movie reviews as containing negative sentiment or not containing negative sentiment. Using one word, does the movie review contain negative sentiment, Yes or No? 0.90
GPT-4 855 Using one word, does the movie review contain negative sentiment, Yes or No? 0.93
GPT-4 855 Using one word, classify the sentiment of the movie review using ’Positive’ or ’Negative’. 0.94
GPT-4 855 You are a researcher who needs to classify movie reviews as containing negative sentiment or not containing negative sentiment. Using one word, does the movie review contain negative sentiment, Yes or No? 0.88
@@
A.3 Analysis of the two tasks using Naïve Bayes classifier with different adjustments
The Naïve Bayes classifier [Kibriya et al., 2005] is a probabilistic algorithm widely used for tasks like text classification due to its simplicity and computational efficiency. It operates on the assumption that features (e.g., words in a document) are conditionally independent given the class label, allowing for straightforward calculation of probabilities. However, the independence assumption is often violated in real-world data, particularly in text classification, where features tend to be highly correlated. This can result in biased probability estimates and reduced classification performance.
Rennie et al. [2003] proposed a solution to address these limitations by introducing TF-IDF weighting and normalization into the Naïve Bayes framework. Their approach adjusts feature importance by emphasizing terms that are frequent in a document but not common across all classes, reducing the classifier’s sensitivity to common words that contribute little to class distinction. Furthermore, they normalized feature vectors to mitigate the influence of document length, ensuring fair comparisons across texts of varying sizes.
Fatourechi et al. [2008] presented another refinement by incorporating logarithmic term frequency weighting. This adjustment reduces the disproportionate impact of extremely high term frequencies, which can dominate classification decisions despite potentially being less informative. By smoothing these term frequencies, their method makes the classifier more robust to outliers and improves its ability to generalize across different datasets.
In this subsection we compare the Naïve Bayes classifiers with the two different fixes described above. We also combine these two into a single classifier, thus having a comprehensive exploration of this method. We end up with four different Naïve Bayes classifiers: the one used in the main body of the manuscript Kibriya et al. [2005] (called Multinomial NB here, which is the same as in the Python package sklearn), the one proposed by Rennie et al. [2003] (called Complement NB), and each one of them with a log scaling as proposed by Fatourechi et al. [2008]. In Figures 13 and 14 we see the F1-macro results for each one of the tasks.
Figure 13: F1 macro scores when performing the Binary Abuse task with the four different Naïve Bayes.
Figure 14: F1 macro scores when performing the Binary Movie Sentiment task with the four different Naïve Bayes.
There is no distinction between the four different classifiers as can be seen in Figures 13 and 14. To understand this, we do two extra analyses. First, in order to understand why the log scaling proposed by Fatourechi et al. [2008] does not seem to have any effect, we perform a z-test between the two feature’s distributions of the vectorized training sets before and after the log scaling. The histogram of the distribution for both tasks can be seen in Figure 15. The z-scores are 0.625 and 0.421 for the binary abuse and binary movie sentiment tasks respectively, showing that the log scaling does not any statistically significant change. Thus, it is expected that we would not observe any difference between the Naïve Bayes classifiers with and without log scaling.
Figure 15: Distribution of features of the vectorized training sets for each task, before and after applying a log scaling.
To understand why there is not any difference between the Multinomial NB classifier [Kibriya et al., 2005] and the complement NB classifier [Rennie et al., 2003], we look at the distribution of weights of the trained Multinomial NB classifier for each label. These can be seen in Figure 16 for the Binary Abuse task and in Figure 17 for the Binary Movie Sentiment task. The four distributions do not show any kind of imbalance, thus making the adjustment of the Complement NB, implemented after Rennie et al. [2003], without effect.
Figure 16: distribution of the weights for each feature of the trained multinomial Naïve Bayes classifier trained for the Binary Abuse task.
Figure 17: distribution of the weights for each feature of the trained multinomial Naïve Bayes classifier trained for the Binary Movie Sentiment task.
A.4 Labelling Functions
@@
Task Type Label Explanation Coverage Overalap
Binary abuse Keywords ABUSE If any of the words considered abusive from a given list of keywords is found in the text, label the text as abusive. 0.36 0.36
Binary abuse Heuristic ABUSE If the length of the text is above 9500 characters (limit of 10 000 characters), label the text as abusive. 0.00 0.00
Binary abuse NLP ABUSE If the polarity of the text is below -0.25, label the text as abusive. 0.07 0.07
Binary abuse NLP NOT_ABUSE If the polarity of the text is equal or more than 0.1, label the text as non abusive. 0.41 0.41
Binary abuse Annotators Both Take the annotators’ data and use it as labelling functions: given an annotator $i$ labelling a set of texts, the annotator-function labels the previously annotated texts and abstains with all non-annotated ones. 0.02 0.02
Binary movie sentiment Keywords Positive If any of the words considered related to positive sentiments is found from a given list of keywords, label the text as positive. 0.90 0.90
Binary movie sentiment Keywords Negative If any of the words considered related to negative sentiments is found from a given list of keywords, label the text as negative. 0.62 0.62
Binary movie sentiment Regex Positive If any expression related to the acting, the direction, the edition, the sound or in general to the film or the movie contains a positive keyword, then label the text as positive. 0.25 0.25
Binary movie sentiment Regex Negative If any expression related to the acting, the direction, the edition, the sound or in general to the film or the movie contains a negative keyword, then label the text as negative. 0.03 0.03
Binary movie sentiment NLP Negative If the polarity of the text is below -0.05, label it as negative. 0.03 0.03
Binary movie sentiment NLP Negative If the polarity of the text is above 0, label it as positive. 0.95 0.95
Binary movie sentiment NLP Positive If the subjectivity of the text is below 0.3, label it as positive. 0.01 0.01
@@
A.5 Zero-shot Prompt Engineering
@@
Task Prompt GPT-3 GPT-3.5 GPT-4 N Non-response
Binary abuse Using one word, does the internet comment contain toxic language, Yes or No? 0.721 0.733 0.781 21
Binary abuse Using one word, is this internet comment using toxic language, Yes or No? 0.667 0.684 0.797 13
Binary abuse You are a researcher who needs to classify comments on the internet as containing abusive language or not containing abusive language. Using one word, does the internet comment contain abusive language, Yes or No? 0.713 0.802 0.813 19
Binary movie sentiment Using one word, does the movie review contain negative sentiment, Yes or No? 0.952 0.923 0.961 2
Binary movie sentiment Using one word, classify the sentiment of the movie review using ’Positive’ or ’Negative’. 0.941 0.945 0.964 8
Binary movie sentiment You are a researcher who needs to classify movie reviews as containing negative sentiment or not containing negative sentiment. Using one word, does the movie review contain negative sentiment, Yes or No? 0.876 0.918 0.939 7
@@
Task Non-response
Binary abuse Cannot determine if the internet comment contains abusive language or not based on the given text
Binary abuse I cannot determine if this internet comment contains abusive language or not as it is not a complete sentence
Binary abuse The internet comment does not contain a clear indication of abusive language
Binary abuse The given internet comment does not contain any language
Binary abuse I cannot determine if the internet comment contains abusive language or not as the comment is incomplete and ends abruptly
Binary movie sentiment Neutral
Binary movie sentiment Mixed